# iWISDM: Assessing instruction following in multimodal models at scale

### v3

@@ 1 Introduction
A typical day in most people’s lives involves hundreds or thousands of tasks. Most of which are performed without explicit attention. Just in between getting up and getting to work, one may have already performed 5-15 tasks (taking a shower, shaving, making coffee, getting dressed, etc.). Teaching artificial agents to perform similar seemingly mundane tasks has proven to be an extremely difficult computational problem (Konar, 2018). The challenge becomes more apparent when one realizes that each of these seemingly mundane tasks such as making coffee involve tens of steps/actions (Figure 1). The challenge becomes increasingly more significant once we consider more complex tasks such as operating a device or assembling a piece of furniture from its instruction manual. And yet, these tasks are performed proficiently by most individuals in most situations.
Large Language Models (LLMs) have become increasingly capable of comprehending natural language across wide topics and contexts, allowing them to hold meaningful conversations, give expert advice, and analyze data among other features (Brown et al., 2020; Ouyang et al., 2022; Radford et al., 2019). In the meantime, their multimodal counterparts are starting to emerge, signalling broader application of such models across industries. Large Multimodal Models (LMMs) are generally capable of receiving and responding in a range of possible modalities including visual, text, and audio (Alayrac et al., 2022; Liu et al., 2023b; Achiam et al., 2023). For example the Gemini-Ultra model, which accepts text, image, audio, and video inputs, and responds with a combination of text and image outputs, recently achieved state of the art on a range of single and multimodal benchmarks (Team et al., 2023).
@@
However, existing benchmarks for assessing such models face several shortcomings: (1) Most multimodal benchmarks like FLEURS (audio-based, (Conneau et al., 2023)) and VATEX (video-based, (Wang et al., 2019)) are still unimodal in their inputs, and do not permit detailed assessment of models’ capacity to integrate information across modalities towards task goals. (2) Visual Question-Answering (VQA) datasets like VQAv2 (Goyal et al., 2017) and CLEVR (Johnson et al., 2017) assess reasoning with visual information in static images without addressing temporal information integration and sequential decision-making. (3) Open-ended learning environments such as XLand (Team et al., 2021), Crafter (Hafner, 2021), and Minecraft (Guss et al., 2019) have been utilized for training reinforcement learning agents, but it remains unclear whether and how they can be adapted to benchmark LMMs. (4) To the best of our knowledge, none of the existing benchmarks specifically assess models’ ability to precisely follow instructions in the context of decision making tasks, an important measure of reliability and trustworthiness. Despite its importance, conducting such assessments have been particularly challenging because of the difficulty of collecting samples of multi-step tasks with ground truth information. (5) More recent benchmarks such as MME (Fu et al., 2023), Mmbench (Liu et al., 2023c) and MMvet (Yu et al., 2023) cover a wide range of cognitive tasks that start to adopt manually-generated or GPT-powered responses. However, those benchmarks are difficult to scale, which makes them inconvenient when investigating the scaling properties of LMMs. In addition, benchmarks such as OwlEval (Ye et al., 2023) and LVLMeHub (Xu et al., 2023) rely on subjective human responses that often show high variability across individuals.
@@
    We introduce iWISDM, a virtual environment for procedural generation of limitless visual decision-making tasks accompanied by natural language instructions.
@@ 3.1 Design
Inspired by prior work (Yang et al., 2018), iWISDM generates tasks through a 3-phase procedure: (1) Task graph construction; (2) Node initialization and; (3) Trial instantiation. Dividing the generation procedure into distinct phases eliminates the necessity of constructing the task graph anew for each task trial. Once all properties associated with the nodes in a given task graph have been specified (phase 2), the user can generate any number of trials of that particular task, each with potentially different stimuli, ground-truth actions, as well as any other inherent stochastic values such as the number of delay frames. In general, iWISDM creates tasks following:
@@ 3.1.1 Task graph construction
To ensure each task graph would comply with the connectivity rules between operators, we follow a backwards initialization process during AutoTask (Figure A1). The task generation process starts from the root node and descends recursively. For each current node/operator $n$ in the graph, its downstream operators $C_{n}$ are randomly sampled based on the connectivity rules. As the graph depth approaches the specified maximum depth, the permissible operators with the shortest possible subtask depth are sampled into $C_{n}$. For instance in our core build, if $n$ is the And operator, then only IsSame, NotSame are sampled since they have shorter subgraph depths than And, Or. Through this procedure, iWISDM AutoTask facilitates sampling from diverse task spaces with varying degrees of complexity specified by the user. The utilization of the connectivity rule dictionary and Switch operator guarantees the logical and feasible nature of the generated tasks, providing researchers with an extensive pool of tasks for investigation and exploration.
@@ 3.1.3 Task trial instantiation
Figure 3: Description of operators and task graph examples in the core build of iWISDM. a) Two main types of operators are considered. Functional Operators: Select operators retrieve stimuli according to specified attributes such as time and location; Switch operators accept boolean values and direct the task logic to one of two different subtasks; GetAttribute operators take stimuli as input and output the corresponding attribute values of those stimuli. Boolean Operators comprise Exit, And, Or, IsSame, and NotSame. These operators take boolean values as inputs and produce another boolean value based on their boolean logic. b-d) Using the operators defined in panel a, we can create tasks following pre-specified rules. Panels b, c, and d display the task graphs for the three typical cognitive tasks illustrated in Figure 2b, demonstrating how these operators are interconnected to define the structures of these specific tasks. For instance, the task instruction for panel b and c are: “category of object 1 equals category of object 2?”, “if category of object 1 equals category of object 3, then category of object 2 equals category of object 3, else category of object 2 equals category of object 4?’. The task instruction for panel c is similar to panel b but repeated across time.
@@ 3.2 Main Features
Compositionality. Real-world tasks are fundamentally compositional, as most tasks can be readily decomposed into sets of simpler subtasks that involve fewer sensory observations and cognitive operations. Two crucial facets of compositionality need to be considered: logical and temporal, both of which are common in daily human behavior and allow individuals to efficiently handle complex tasks and adapt to dynamic environments. Cognitive processes involving task decomposition (i.e. breaking tasks into subtasks) and temporal combination of decision rules (i.e. combining the outcomes of subtasks) are fundamental to our ability to navigate the world around us.
@@ A.2 Additional Figures
Figure A5: An example trial with distractors. Instruction: ”observe object 1, observe object 2, category of object 2 not equals category of object 1? observe object 3,observe object 4 with location:, top left, category of object 4 equals couches or identity of object 3 equals identity of object 1? delay, observe object 5, observe object 6, delay, observe object 7, identity of object 7 equals identity of object 6 and identity of object 4 equals identity of object 3?”
@@
Figure A7: Additional Analyses of Open-Source model results (xcomposer2 & MMICL)

### v4

@@ 1 Introduction
A typical day in most people’s lives involves hundreds or thousands of tasks. Most of which are performed without explicit attention. Just in between getting up and getting to work, one may have already performed 5-15 tasks (taking a shower, shaving, making coffee, getting dressed, etc.). Teaching artificial agents to perform similar seemingly mundane tasks has proven to be an extremely difficult computational problem (Konar, 2018). The challenge becomes more apparent when one realizes that each of these seemingly mundane tasks such as making coffee involves tens of steps/actions (Figure 1). The challenge becomes increasingly more significant once we consider more complex tasks such as operating a device or assembling a piece of furniture from its instruction manual. And yet, these tasks are performed proficiently by most individuals in most situations.
Large Language Models (LLMs) have become increasingly capable of comprehending natural language across wide topics and contexts, allowing them to hold meaningful conversations, give expert advice, and analyze data among other features (Brown et al., 2020; Ouyang et al., 2022; Radford et al., 2019). In the meantime, their multimodal counterparts are starting to emerge, signalling broader application of such models across industries. Large Multimodal Models (LMMs) are generally capable of receiving and responding in a range of possible modalities including visual, text, and audio (Alayrac et al., 2022; Liu et al., 2023b; Achiam et al., 2023). For example, the Gemini-Ultra model, which accepts text, image, audio, and video inputs, and responds with a combination of text and image outputs, recently achieved state-of-the-art on a range of single and multimodal benchmarks (Team et al., 2023).
@@
However, existing benchmarks for assessing such models face several shortcomings: (1) Most multimodal benchmarks like FLEURS (audio-based, (Conneau et al., 2023)) and VATEX (video-based, (Wang et al., 2019)) are still unimodal in their inputs, and do not permit detailed assessment of models’ capacity to integrate information across modalities towards task goals. (2) Visual Question-Answering (VQA) datasets like VQAv2 (Goyal et al., 2017) and CLEVR (Johnson et al., 2017) assess reasoning with visual information in static images without addressing temporal information integration and sequential decision-making. (3) Open-ended learning environments such as XLand (Team et al., 2021), Crafter (Hafner, 2021), and Minecraft (Guss et al., 2019) have been utilized for training reinforcement learning agents. It remains unclear whether and how they can be adapted to benchmark LMMs. (4) To the best of our knowledge, none of the existing benchmarks specifically assess models’ ability to precisely follow instructions in the context of decision making tasks, an important measure of reliability and trustworthiness. Despite its importance, conducting such assessments has been particularly challenging because of the difficulty of collecting samples of multi-step tasks with ground truth information. (5) More recent benchmarks such as MME (Fu et al., 2023), Mmbench (Liu et al., 2023c) and MMvet (Yu et al., 2023) cover a wide range of cognitive tasks that start to adopt manually-generated or GPT-powered responses. However, those benchmarks are difficult to scale, which makes them inconvenient when investigating the scaling properties of LMMs. In addition, benchmarks such as OwlEval (Ye et al., 2023) and LVLMeHub (Xu et al., 2023) rely on subjective human responses that often show high variability across individuals.
@@
    We introduce iWISDM, a virtual environment for the procedural generation of limitless visual decision-making tasks accompanied by natural language instructions.
@@ 3.1 Design
Inspired by prior work (Yang et al., 2018), iWISDM generates tasks through a 3-phase procedure: (1) Task graph construction; (2) Node initialization (3) Trial instantiation. Dividing the generation procedure into distinct phases eliminates the necessity of constructing the task graph anew for each task trial. Once all properties associated with the nodes in a given task graph have been specified (phase 2), the user can generate any number of trials of that particular task, each with potentially different stimuli, ground-truth actions, as well as any other inherent stochastic values such as the number of delay frames. In general, iWISDM creates tasks following:
@@ 3.1.1 Task graph construction
To ensure each task graph would comply with the connectivity rules between operators, we follow a backward initialization process during AutoTask (Figure A1). The task generation process starts from the root node and descends recursively. For each current node/operator $n$ in the graph, its downstream operators $C_{n}$ are randomly sampled based on the connectivity rules. As the graph depth approaches the specified maximum depth, the permissible operators with the shortest possible subtask depth are sampled into $C_{n}$. For instance in our core build, if $n$ is the And operator, then only IsSame, NotSame are sampled since they have shorter subgraph depths than And, Or. Through this procedure, iWISDM AutoTask facilitates sampling from diverse task spaces with varying degrees of complexity specified by the user. The utilization of the connectivity rule dictionary and Switch operator guarantees the logical and feasible nature of the generated tasks, providing researchers with an extensive pool of tasks for investigation and exploration.
@@ 3.1.3 Task trial instantiation
Figure 3: Description of operators and task graph examples in the core build of iWISDM. a) Two main types of operators are considered. Functional Operators: Select operators retrieve stimuli according to specified attributes such as time and location; Switch operators accept boolean values and direct the task logic to one of two different subtasks; GetAttribute operators take stimuli as input and output the corresponding attribute values of those stimuli. Boolean Operators comprise Exit, And, Or, IsSame, and NotSame. These operators take boolean values as inputs and produce another boolean value based on their boolean logic. b-d) Using the operators defined in panel a, we create tasks from pre-specified rules. Panels b, c, and d display the task graphs for the three typical cognitive tasks illustrated in Figure 2b, demonstrating how these operators are interconnected to define the structures of these specific tasks. For instance, the task instruction for panel b and d are: “category of object 1 equals category of object 2?”, “if category of object 1 equals category of object 3, then category of object 2 equals category of object 3, else category of object 2 equals category of object 4?’. The task instruction for panel c is similar to panel b but repeated across time.
@@ 3.2 Main Features
Compositionality. Real-world tasks are fundamentally compositional, as most tasks can be readily decomposed into sets of simpler subtasks that involve fewer sensory observations and cognitive operations. Two crucial facets of compositionality need to be considered: logical and temporal, both of which are common in daily human behaviour and allow individuals to efficiently handle complex tasks and adapt to dynamic environments. Cognitive processes involving task decomposition (i.e. breaking tasks into subtasks) and temporal combination of decision rules (i.e. combining the outcomes of subtasks) are fundamental to our ability to navigate the world around us.
@@ A.2 Additional Figures
Figure A5: An example trial with distractors. Instruction: ”observe object 1, observe object 2, category of object 2 not equals category of object 1? observe object 3,observe object 4 with location: top left, category of object 4 equals couches or identity of object 3 equals identity of object 1? delay, observe object 5, observe object 6, delay, observe object 7, identity of object 7 equals identity of object 6 and identity of object 4 equals identity of object 3?”
@@
Figure A7: Average accuracy of GPT-4v, xcomposer2, and MMICL across a varying number of delay frames.